### v2

# Pierce the Mists, Greet the Sky: Decipher Knowledge Overshadowing via Knowledge Circuit Analysis

Abstract
Large Language Models (LLMs), despite their remarkable capabilities, are hampered by hallucinations. A particularly challenging variant, knowledge overshadowing, occurs when one piece of activated knowledge inadvertently masks another relevant piece, leading to erroneous outputs even with high-quality training data. Current understanding of overshadowing is largely confined to inference-time observations, lacking deep insights into its origins and internal mechanisms during model training. Therefore, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect knowledge overshadowing. By innovatively employing knowledge circuit analysis, PhantomCircuit dissects the internal workings of attention heads, tracing how competing knowledge pathways contribute to the overshadowing phenomenon and its evolution throughout the training process. Extensive experiments demonstrate PhantomCircuit ’s effectiveness in identifying such instances, offering novel insights into this elusive hallucination and providing the research community with a new methodological lens for its potential mitigation.
[…]
tcb@breakable
Pierce the Mists, Greet the Sky:
Decipher Knowledge Overshadowing via Knowledge Circuit Analysis
Haoming Huang1,3, [^1] Co-first authors with equal contribution., Yibo Yan1,2, [^1] Co-first authors with equal contribution., Jiahao Huo1, Xin Zou1, Xinfeng Li4, Kun Wang4, Xuming Hu1,2, [^2] Corresponding author. 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology 3Shanghai Jiao Tong University   4Nanyang Technological University
1 Introduction
Large Language Models (LLMs) have witnessed explosive growth in recent years, demonstrating remarkable capabilities across a multitude of domains, including natural language understanding, generation, reasoning, and even cross-modal tasks Chang et al. (2024); Zhao et al. (2024); Yan et al. (2025a, 2024a); Xun et al. (2025). Their proficiency has catalyzed transformative advancements in various applications. However, a persistent challenge that tempers their widespread adoption and reliability is the phenomenon of hallucination. Broadly, hallucinations refer to instances where models generate content that is factually incorrect, nonsensical, or unfaithful to the provided source context, despite appearing coherent and fluent Rawte et al. (2023); Chakraborty et al. (2025); Wang et al. (2025).
Figure 1: Illustrative comparison of previous research with inference-time analysis (b) and observation-based explanation (c) vs our proposed PhantomCircuit (d) on knowledge overshadowing (a).
While substantial research has been dedicated to understanding the causes and detection of general hallucinations, a specific variant known as “knowledge overshadowing” warrants deeper investigation Zhang et al. (2024b, 2025a). This phenomenon is particularly perplexing because it can manifest even when models are trained on high-quality, meticulously curated pre-training corpora. Current understanding, primarily derived from inference-time observations, characterizes overshadowing as a scenario where, for a given query, one piece of activated knowledge inadvertently “overshadows” another relevant knowledge. This interference ultimately biases the model’s reasoning process, leading to a hallucinatory output, as illustrated in Figure 1 (a).
Nevertheless, existing explorations into knowledge overshadowing suffer from notable limitations. ❶ They predominantly focus on inference-time analysis, as shown in Figure 1 (b). While valuable for identifying the occurrence of overshadowing, such observations offer a surface-level understanding and often fall short of elucidating how these detrimental patterns are learned during the training phase. ❷ The explanations for overshadowing are often speculatively inferred from these observational outcomes rather than being rigorously investigated through dedicated interpretability tools that can probe the model’s internal decision-making mechanisms, as shown in Figure 1 (c). Consequently, a more comprehensive analytical framework is imperative to dissect this phenomenon from its origins to its manifestation.
To bridge this gap, we introduce PhantomCircuit, a novel framework designed to comprehensively analyze and detect the knowledge overshadowing phenomenon. Specifically, PhantomCircuit facilitates an in-depth examination of the evolution of overshadowing hallucinations throughout the training process, correlating their emergence and prevalence with core factors such as knowledge popularity, model size, and dataset size. Then, by leveraging knowledge circuit analysis as a key interpretability technique, we aim to trace the flow of information and the formation of knowledge representations within attention heads, thereby uncovering the internal mechanisms giving rise to overshadowing. Furthermore, we propose to optimize the number of edges within these circuits, thus alleviating the knowledge overshadowing. As illustrated in Figure 1 (d), our overall work aims to provide a clearer, mechanistic understanding and potential strategies for this elusive type of hallucination.
Our contributions can be summarized as follows:
  • We introduce PhantomCircuit, the first comprehensive framework designed to systematically analyze and detect knowledge overshadowing, delving into its mechanistic nature and evolution throughout model training.
  • We pioneer the use of knowledge circuit analysis to dissect the internal workings of attention heads, specifically elucidating how competing knowledge pathways contribute to the overshadowing phenomenon.
  • We conduct extensive experiments to demonstrate PhantomCircuit ’s efficacy in detecting knowledge overshadowing, offering novel insights and a new methodological lens for the research community.
2 Related Work
2.1 Hallucination Detection
Factuality hallucination detection, which aims to evaluate whether the output of LLMs aligns with real-world facts, typically involves either external fact-checking or internal uncertainty analysis Huang et al. (2025); Dang et al. (2024); Zheng et al. (2024); Zou et al. (2024); Zhou et al. (2024); Zhu et al. (2024). For instance, FACTSCORE Min et al. (2023) decomposes a generation into atomic facts and calculates the proportion that are supported by reliable knowledge sources. FACTOOL Chern et al. (2023), on the other hand, integrates multiple tools such as Google Search and Google Scholar to gather external evidence and assess the factuality of generated content. In contrast, methods like Chain-of-Verification Dhuliawala et al. (2023), probability-based assessments Kadavath et al. (2022); Zhang et al. (2024a), and uncertainty estimation approaches Varshney et al. (2023); Yao et al. (2023); Luo et al. (2023) rely on LLMs’ internal parametric knowledge or uncertainty signals to predict potential hallucinations. Among these efforts, knowledge overshadowing Zhang et al. (2025a) offers a novel perspective by modeling hallucination behavior from the perspective of knowledge representation, providing an efficient strategy for proactive prevention.
2.2 Knowledge Circuit Analysis
In the context of mechanistic interpretability Rai et al. (2024); Huo et al. (2024); Huang et al. (2024a), computations in Transformer-based language models are viewed as a connected directed acyclic graph encompassing components such as MLPs and attention layers Syed et al. (2023); Conmy et al. (2023); Huang et al. (2024b). A circuit refers to a sparse computational subgraph that significantly influences the model’s behavior on a specific task Olah et al. (2020); Elhage et al. (2021); Wang et al. (2022). Building on this, Yao et al. (2024) introduce the concept of knowledge circuits, hypothesizing that cooperation among components reveals implicit knowledge in LLMs. Further, Ou et al. (2025) explore how such circuits evolve during continual pre-training, providing insights into knowledge acquisition. To enable effective knowledge editing, CaKE Yao et al. (2024) proposes a Circuit-aware Knowledge Editing method that guides models to activate modified knowledge and form new reasoning pathways. In this paper, we analyze the phenomenon of knowledge overshadowing through the lens of knowledge circuits, contributing new perspectives to LLM hallucination detection.
See more related work in Appendix A.1.
3 Methodology
This section first defines the knowledge overshadowing phenomenon and its quantitative evaluation. Subsequently, we detail the PhantomCircuit framework, which encompasses methods for analyzing its training dynamics and for constructing and analyzing knowledge circuits to understand its internal mechanisms [^1] Our code will be available upon acceptance..
3.1 Knowledge Overshadowing
Knowledge overshadowing refers to a specific type of hallucination where less prevalent, subordinate knowledge is suppressed by high-frequency, dominant knowledge when both are associated with common background knowledge Zhang et al. (2025a).
Let $X_{dom}$ denote dominant knowledge entities and $X_{sub}$ denote subordinate knowledge entities, both potentially co-occurring with background knowledge $X_{bg}$. The core idea is that a strong learned pattern $X_{dom}\leftrightarrow X_{bg}$ can "over-generalize" where the model primarily associates $X_{dom}$ with $X_{bg}$.
Consequently, when the model encounters $X_{sub}$ with $X_{bg}$, denoted as $X_{sub}\leftrightarrow X_{bg}$), it may erroneously favor outputs related to $X_{dom}$ due to the stronger $X_{dom}\leftrightarrow X_{bg}$ pattern.
3.1.1 Knowledge Overshadowing Occurrence
When the input prompt is a knowledge pair composed of background knowledge and dominant knowledge, denoted as $P_{dom}=(X_{bg},X_{dom})$, and the model correctly generates the answer corresponding to the dominant knowledge, denoted as $Y_{dom}$, we consider this outcome, represented by the pair $(P_{dom},Y_{dom})$, as a successful recall of dominant knowledge.
When the input prompt is $P_{sub}$, but the model wrongly generates $Y_{dom}$, knowledge overshadowing occurs for this query-response instance, resulting in the $(P_{sub},Y_{dom})$.
3.1.2 Quantitative Indicators
Let $N_{dom}$ and $N_{sub}$ be the number of instances of $P_{dom}$ and $P_{sub}$ in the training dataset, respectively. The dataset $Z$ comprises a subset $Z_{dom}$ containing $N_{dom}$ instances of $P_{dom}$, and a subset $Z_{sub}$ containing $N_{sub}$ instances of $P_{sub}$.
During autoregressive generation tasks performed by the model, let $M_{sub}$ be the number of times when overshadowing instances $(P_{sub},Y_{dom})$ occur , and $M_{dom}$ be the number of times $(P_{dom},Y_{dom})$ occurs. Then, we can define the absolute extent of the knowledge overshadowing effect, the Absolute Overshadowing rate ($\mathcal{AO}$) and calculate it using $M_{sub}$ and $N_{sub}$,
$\mathcal{AO}=p(Y_{dom}|P_{sub})=\frac{M_{sub}}{N_{sub}}.$ (1)
To account for the model’s inherent performance and potential noise affecting the overshadowing rate, we also introduce $R_{dom}$ for dominant knowledge inputs:
$R_{dom}=p(Y_{dom}|P_{dom})=\frac{M_{dom}}{N_{dom}},$ (2)
which represents the recall rate for $P_{dom}$ query-response instances.
The Relative Overshadowing rate ($\mathcal{RO}$) is then defined as:
$\mathcal{RO}=\frac{\mathcal{AO}}{R_{dom}}=\frac{p(Y_{dom}|P_{sub})}{p(Y_{dom}|P_{dom})}.$ (3)
3.1.3 Overshadowing Influence Factors
Knowledge Popularity (P) is the fundamental cause of the knowledge overshadowing phenomenon and serves as its primary influencing factor. P is defined as the ratio of the capacities of $Z_{dom}$ to $Z_{sub}$, thus $P=N_{dom}/N_{sub}.$
Model Size (M), referring to the number of model parameters, also impacts knowledge overshadowing. A larger M generally implies stronger generalization capabilities, causing the model to rapidly generalize the $X_{dom}\leftrightarrow X_{bg}$ to instances involving $X_{sub}\leftrightarrow X_{bg}$ and exacerbating overshadowing.
In addition to the factors mentioned in Zhang et al. (2025a, 2024b), aiming to analyze the dynamic evolution of knowledge overshadowing during the training process, we extend our consideration to total number of tokens, the Dataset Size (D) in the training set. The average loss proportion of subordinate knowledge $\mathcal{LP}$ within an epoch is defined as $\mathcal{LP}=loss(P_{sub})/total\;loss,$ which also relates to the overshadowing.
3.2 Analysis Framework PhantomCircuit
Our proposed knowledge circuit-based overshadowing analysis framework involves the overshadowing dynamics analysis during training and circuit-based internal mechanism analysis.
3.2.1 Overshadowing Dynamic Analysis
Our framework provides a novel dynamic analysis of knowledge overshadowing during model training. By manipulating P, M, and D, we monitor $\mathcal{RO}$ across epochs. We focus on identifying the onset, duration, and recovery stages of overshadowing to understand their modulation by P, M, and D. Recognizing P as a key factor, we also explore how $\mathcal{LP}$ co-evolves with $\mathcal{RO}$ under these variables, aiming to uncover the role of $\mathcal{LP}$ in explaining overshadowing dynamics.
3.2.2 Knowledge Circuit Construction
We construct the knowledge circuit, a sparse computational subgraph $C\subseteq G$, where $G=(V,E)$ is the directed acyclic graph representation of LLMs, $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, $E$ represents the information flow between these components. Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of given input prompt, particularly in knowledge overshadowing, is $\{X_{dom},X_{sub}\}$, the difference between $P_{dom}$ and $P_{sub}$. The adapted construction methods is similar to edge attribution patching Conmy et al. (2023), which involves:
Paired inputs. For a given $X_{bg}$, we create a pair of input prompts, the clean input $P_{sub}$ corresponding to our expected output $Y_{sub}$ in the overshadowing case, and the corrupt input $P_{dom}$ serving for the contrasting mentioned below.
Activation difference calculation. After running this pair of inputs through the model, we calculate the difference in activation values $\Delta A(v_{p}),\Delta A(v_{c})$ for the parent node $v_{p}$ and child node $v_{c}$ of each edge under these distinct inputs, $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corrupt}(v_{p}),$ where $A_{clean}(v_{p})$ and $A_{corrupt}(v_{p})$ are the activation of parent node when $P_{sub}$ as clean input and $P_{dom}$ as corrupt input.
Edge score calculation. The importance $S(e)$ of an edge $e=(v_{p},v_{c})$ is scored based on how patching $v_{p}$’s activation (using $\Delta A(v_{p})$) influences a metric $\mathcal{M}$, which assesses the model’s ability to correctly output $Y_{sub}$ rather than $Y_{dom}$ for $P_{sub}$ inputs. Using the Integrated Gradients, $S(e)$ is approximated as:
$S(e)\approx Exp\left[\Delta A(v_{p})\cdot\frac{\partial\mathcal{M}(Y_{sub}|P_{sub})}{\partial A(v_{p})}\right].$ (4)
Circuit construction. Edges with scores $|S(e)|$ below a threshold $\tau$ are pruned. The remaining subgraph forms the knowledge circuit $C$. See more details about construction in Appendix B.1
3.2.3 Circuit-based Analysis
PhantomCircuit mainly focuses on the attention heads in $C$ and follows these steps:
Node Attention Analysis. We identify high attention heads within $C$ by examining their attention scores and patterns, specifically their focus on $\{X_{dom},X_{sub}\}$.
Circuit Structure Analysis. We then trace the information flow of these high attention heads by identifying their parent and child nodes to understand their structural role. Nodes consistently retained in circuits built with different thresholds $\tau$ are also analyzed as key components.
Layer-wise Logit Evolution. Finally, using logit lens nostalgebraist (2020), we inspect the evolving output logits at layers associated with key nodes. This validates if their captured information contributes to the model’s prediction as expected.
3.2.4 Circuit-based Overshadowing Recovery
Inspired by the goal of knowledge circuits to maximize sensitivity to the distinguishing features between $X_{dom}$ and $X_{sub}$, we propose a circuit-based method to alleviate overshadowing. This involves optimizing the circuit structure by tuning the edge pruning threshold $\tau$ to obtain an optimal circuit, $C_{opt}$. The optimization is formulated as:
$\tau_{opt}=\operatorname*{arg\,max}_{\tau}\mathcal{M}(C_{opt}(\tau),P_{sub},Y_{sub}),$ (5)
where $\mathcal{M}$ measures the circuit’s ability to distinguish $\{X_{dom},X_{sub}\}$, This results in a $C_{opt}(\tau_{opt})$ which is expected to mitigate or eliminate the overshadowing effect for specific input prompts.
To automate the overshadowing recovery process and extend its applicability, we simplify the relative pointwise mutual information (R-PMI) method from Zhang et al. (2025a, 2024b) to identify $X_{sub}$ within $P_{sub}$. First, we obtain the top-$k$ next-token candidates, $V_{top}(P_{sub})$, by feeding $P_{sub}$ to the model. Then, we iteratively generate contrastive prompts $P^{\prime}_{sub}$ by masking (in our implementation, by deleting) each token $X^{\prime}_{sub}$ from $P_{sub}$. For each $P^{\prime}_{sub}$, we acquire its top-$k$ candidates $V_{top}(P^{\prime}_{sub})$. The R-PMI for each token $y_{i}$ in the intersection $V_{top}(P_{sub})\cap V_{top}(P^{\prime}_{sub})$ is calculated as:
$R\textnormal{-}PMI_{i}=\log\frac{p(y_{i}|P_{sub})}{p(y_{i}|P^{\prime}_{sub})}.$ (6)
Then, we sum the negative R-PMI values to get $S_{R\textnormal{-}PMI}=\sum\min(R\textnormal{-}PMI_{i},0)$. The $X^{\prime}_{sub}$ yielding the minimum $S_{R\textnormal{-}PMI}$ is identified as $X_{sub}$.
Furthermore, the $Y_{sub}$ is determined as the token $y_{i}$ from $V_{top}(P_{sub})$ whose average rank improves most significantly when non-subordinate components $X^{\prime}_{sub}$, often the $X_{bg}$, is masked. $Y_{dom}$ is identified as $y_{i}$ that has the highest average rank across all $P^{\prime}_{sub}$. For circuit construction, $X_{dom}$ within $P_{dom}$ is replaced by a generic placeholder like "something", as shown in Figure 5. Combining this streamlined approach enables broader application of our knowledge circuit-based overshadowing recovery. See more details in Appendix B.2
(a) Popularity (P)
(b) Model Size (M)
(c) Dataset Size (D)
(d) Popularity (P)
(e) Model Size (M)
(f) Dataset Size (D)
Figure 2: (a) $\sim$ (c) show the dynamic variation of $\mathcal{RO}$ relating to P, M and D in model training phase. Higher Knowledge Popularity (P) and Model Size (M) tend to result in an earlier onset, shorter duration, and quicker recovery from knowledge overshadowing. In contrast, a larger Dataset Size (D) also leads to an earlier onset but is associated with a slower recovery phase. (d) $\sim$ (f) show the duration of onset stage, high $\mathcal{RO}$ (> 90%) stage as well as recovery stage, and $\mathcal{RO}$’s rate of change during the onset and recovery stages.
4 Experiment
4.1 Experiment Setup
(a) M = 70M
(b) M = 410M
(c) M = 1.4B
Figure 3: The co-evolution between $\mathcal{RO}$ $\mathcal{LP}$ in different M. Early high overall loss and low $\mathcal{LP}$ leads to intensive optimization of $P_{dom}$ and $\mathcal{RO}$ rises up to near 100%. As training progresses, subordinate knowledge loss proportion ($\mathcal{LP}$) rises, shifting optimization focus to $P_{sub}$ errors, initiating $\mathcal{RO}$’s recovery phase, validated across models.
4.1.1 Dataset
Synthetic dataset. To investigate the dynamic characteristics and influencing factors of the $\mathcal{RO}$ during training under controlled conditions,and to minimize the complexities and semantic relationships inherent in natural language, we construct a synthetic dataset and train models from scratch.
Follow Zhang et al. (2025a), we first fix the length of $X_{bg}$ at 4 tokens and the lengths of $X_{dom},X_{sub},Y_{dom}$, $Y_{sub}$ at 1. For the dataset size (D), we experiment with 0.26 (D = 0.26M), 2.6 (D = 2.6M) and 26 million tokens (D = 26M). For popularity (P), we set P values as 5, 25, and 100.
Then, for a specific combination of P and D, there are several distinct groups to achieve the target D. Each group comprises P distinct $X_{dom}$ and a single $Y_{dom}$ for dominant knowledge, one $X_{sub}$ and $Y_{sub}$ for subordinate knowledge. The $X_{bg}$ is shared. Thus, there are P unique $\{P_{dom},Y_{dom}\}$ and one $\{P_{sub},Y_{sub}\}$ in a group. All the tokens are randomly sampled. See more details of our dataset in Appendix C.
Finetuning Dataset. We construct a finetuning dataset to evaluate circuit-base overshadowing recovery method. Utilizing virtual knowledge, we preserve the natural language semantics while avoiding prior knowledge editing that could interfere with the natural occurrence of overshadowing. For this dataset, we set P = 5 and D = 1M. The Appendix D shows the dynamics analysis based on finetuning dataset.
4.1.2 Models Evaluated
We employ models from the Pythia suite Biderman et al. (2023), specifically: Pythia-70M, Pythia-410M, Pythia-1.4B, and Pythia-2.8B, corresponding to model sizes (M) of M = 70M, 410M, 1.4B and 2.8B, respectively. Tokens randomly sampled to build synthetic dataset is from Pythia tokenizer.
4.1.3 Training
A uniform learning rate of $10^{-5}$ and batch size of 16 are used for both dataset. Training is conducted on NVIDIA A800 GPUs.
4.1.4 Evaluation
To measure $\mathcal{RO}$, we randomly sample 500 $P_{dom}$ and 500 $P_{sub}$ prompts for evaluation after each training epoch. The $\mathcal{LP}$ is recorded within each epoch. The results are shown in Figures 2 and 3.
Figure 4: The circuit-based analysis of proposed PhantomCircuit. (a) shows the average attention scores allocated to $\{X_{dom},X_{sub}\}$ across epoch. (b) focuses on the onset and recovery phases and shows that the appear and disappear of high attention heads (attention score greater than 0.2) attribute to the reduce and raise of $\mathcal{RO}$. (C) shows the logits and ranks of $Y_{sub},Y_{dom}$ across layers. The main structure of circuit (with 400 edges totally) and attention mode show that high attention head a5.h1 and MLP layer m5 contribute to the juncture of $Y_{sub}$ ’s rank. Solid lines denote direct information flow, while dashed lines indicate indirect flow in circuit structure map.
4.2 Main Result
Based on the experiments described above, using the circuit to analyze and optimize overshadowing, our investigation yields the following findings.
A higher value of P and M can lead to an earlier onset, shorter duration, and quicker recovery of the knowledge overshadowing. Distinctly, a larger D contributes to the earlier onset but also a slower recovery from overshadowing.
As shown in Figures 2(a) and 2(d) (M=70M, D=2.6M), increasing P (from 5 to 100) significantly shortens or even eliminates the onset phase. This is attributed to more prominent dominant patterns $X_{dom}\leftrightarrow X_{bg}$ being learned and generalized rapidly, even within the first epoch. The duration phase also decreases with higher P, because a larger P and a fixed D implies fewer groups of knowledge pairs, leading to less diversity of $P_{sub}$, allowing the model to learn all overshadowed $P_{sub}$ and recover from overshadowing more quickly.
Figures 2(b) and 2(e) (P=5, D=2.6M) demonstrate that larger models (M) exhibit shorter or absent onset phases, indicating an earlier occurrence of knowledge overshadowing. This is due to the stronger generalization capabilities of larger models, leading them to quickly learn and overgeneralize dominant patterns. However, larger models also show a shortened duration phase and a rapid decline in $\mathcal{RO}$ during recovery, suggesting enhanced capacity to differentiate $\{X_{dom},X_{sub}\}$, thus recovering faster despite earlier overshadowing.
In Figures 2(c) and 2(f) (P=5, M=70M) larger D leads to an earlier onset of overshadowing, with $\mathcal{RO}$ peaking sooner. This is because more data per epoch provides more iterations and exposure to the dominant pattern, accelerating its generalization. Conversely, the recovery phase is prolonged with larger datasets. The increased diversity of $P_{sub}$ in larger datasets requires more epochs for the model to learn all instances and recover.
Figure 5: Overshadowing recovery via optimized circuit. In the finetuned model, $X_{dom}\leftrightarrow X_{bg}$ causes the original full model to incorrectly predict $Y_{dom}$. First, we detect the $X_{sub},Y_{sub},Y_{dom}$ automatically by calculating the minimum $S_{R\textnormal{-}PMI}$ Then, optimizing the knowledge circuit by pruning edges to keep only key nodes enhances the attention on $\{X_{dom},X_{sub}\}$, enabling the recovery from overshadowing and facilitating correct $Y_{sub}$ prediction.
Notably, across various parameter combinations, $\mathcal{RO}$ often approaches 100% in the early training stages and finally recovers to 0%. We hypothesize this phenomenon stems from initial high-loss state, where optimization efforts disproportionately focus on reducing the larger loss contribution from $P_{dom}$. Our subsequent observations regarding $\mathcal{LP}$ corroborate this.
The dynamic nature of knowledge overshadowing arises from the co-evolution relationship between the $\mathcal{LP}$ and $\mathcal{RO}$. As depicted in Figure 3, in the early training stages, when the overall loss is high and the contribution from $P_{sub}$ is small, the optimization process tends to concentrate on the $P_{dom}$ and $\mathcal{RO}$ rapidly approaches 100%. However, as training progresses and $\mathcal{LP}$ begins to rise, eventually nearing its peak, $P_{sub}$ takes a substantial portion of the remaining loss. At this juncture, the model’s optimization efforts shift to focus on these errors from $P_{sub}$. This shift initiates the recovery phase, leading to a decline in $\mathcal{RO}$. Therefore, the insufficient optimization results in the overshadowing, which is consistent across varying M.
The occurrence of overshadowing in pretrained LLMs can be understood with dynamics analysis results above. Initially, large M and D promote the rapid generalization of $X_{dom}\leftrightarrow X_{bg}$ and overshadowing onset. The subsequent inadequate optimization of the $P_{sub}$ is exacerbated by the sheer scale and diversity of training data, which brings prolonged overshadowing effect and exceeded sharp recovery effect from large M (e.g., trillion tokens for training LlaMa-2-7B ).
The knowledge circuit’s attentional allocation to differences between dominant and overshadowed knowledge inputs dictates the extent of knowledge overshadowing. Figure 4 (a) shows the variation of circuit’s average attention scores on $\{X_{dom},X_{sub}\}$ throughout the training phase. The higher average attention alleviates overshadowing, while lower attention exacerbates it.
Figure 4 (b) shows the circuit dynamics across the onset and recovery phases of knowledge overshadowing. These bar plots show the attention scores of individual attention heads in a specific epoch, indicating that when the $\mathcal{RO}$ declines, some attention heads, defined as high attention heads, exhibiting high focus on $\{X_{dom},X_{sub}\}$ emerge within the circuit. The threshold for high attention is set at 0.2 for the length of $X_{sub}$ is one fifth of the length of input prompt in synthetic dataset. Conversely, when knowledge overshadowing intensifies, a subset of these critical attention heads tends to disappear from the circuit.
Furthermore, by focusing on the circuit’s internal mechanisms and structure within a specific epoch, as shown in Figure 4 (c), we leverage the circuit-based analysis of our proposed PhantomCircuit. First, according to the logits and ranks of $Y_{dom}$ and $Y_{sub}$ across layers, the 5th layer is a juncture where the rank of $Y_{sub}$ exceeds $Y_{dom}$. Subsequently, by focusing on the circuit structure, we identify the internal mechanisms driving this juncture. A high attention head, a5.h1 (layer 5), crucially channels information to the subsequent MLP layer, m5. The a5.h1 also appears to inherit its high attention state from earlier layers, mediated by the layer 4 MLP (m4), leading it to not only continue elevating the logits for $Y_{sub}$ but also to attenuate the previously rapid growth of $Y_{dom}$’s logits, thereby facilitating the observed rank reversal. The attention map on the right confirms its high focus on $\{X_{dom},X_{sub}\}$.
Knowledge circuit-guided optimization represents a promising strategy to mitigate the knowledge overshadowing effect. Figure 5 illustrates our findings. We first finetune the model on finetuning dataset. During the recovery phase, we randomly choose some $P_{sub}$. Feeding $P_{sub}$ to the original full model (M = 410M), the prediction is still $Y_{dom}$ instead of expected $Y_{sub}$, and shows a significant gap in probability as well.
We then detect the $Y_{sub},Y_{dom}$ and $X_{sub}$ for chosen $P_{sub}$ automatically, replace the $X_{sub}$ of placeholder token "something". With these components, we obtain the $P_{dom}$ as corrupt input and $P_{sub}$ as clean input to construct a knowledge circuit. A golden section search algorithm is employed to determine the optimal number of edges for building $C_{opt}$. The optimized circuit structure map shows that some attention heads are pruned, which are often low attention heads, or exhibit no significant attentional pattern towards the core task-relevant information. Retained high attention heads are key to differentiating $P_{dom}$ from $P_{sub}$ which give significant attention to $\{X_{dom},X_{sub}\}$. Some low attention heads also remain, implying that even in the circuit, processing background knowledge $X_{bg}$ and linking it to the distinctive elements $X_{sub}$ is crucial for correct inference. The performance of the optimized circuit $C_{opt}$ is then evaluated by feeding it the clean input $P_{sub}$, while $P_{dom}$ serves as the baseline for contrast. Finally, the circuit successfully produces $Y_{sub}$, demonstrating the elimination of the overshadowing effect.
Future work will enhance the circuit optimization metric $\mathcal{M}$ by incorporating $Y_{sub}$’s absolute logit alongside the logit difference with $Y_{dom}$ for more effective guidance. Developing a comprehensive evaluation framework for circuit-based recovery is also crucial. These steps will evolve PhantomCircuit into an integrated platform for efficient analysis and robust optimization of knowledge overshadowing.
5 Conclusion
This paper investigates hallucinations in LLMs caused by knowledge overshadowing, and introduces PhantomCircuit, a novel knowledge circuit-based analysis framework. PhantomCircuit first analyzes the training dynamics of overshadowing, finding that dominant knowledge popularity, model size, and dataset size critically shape the onset, duration, and recovery of overshadowing. Apart from that, the persistent overshadowing in pretrained models stems from inadequately optimized subordinate knowledge loss. By analyzing knowledge circuits, we find that changes in critical attention heads’ focus on subordinate knowledge directly correlate with the recovery or onset of overshadowing. Finally, optimizing these knowledge circuits presents a promising strategy for mitigating knowledge overshadowing.
Limitations
Despite the insights provided by PhantomCircuit, this study has several limitations that open avenues for future research:
  1. 1.
    The dynamic analysis of knowledge circuits throughout training is computationally intensive, potentially hindering scalability to very large models or extensive training. We aim to develop more computationally efficient techniques for approximating circuit evolution, such as checkpoint-based analysis or lightweight probing.
  2. 2.
    This study concentrates on a specific type of knowledge overshadowing, leaving more complex or subtle interference patterns unaddressed. Future work will broaden PhantomCircuit’s scope to investigate a wider range of overshadowing phenomena, including those in multi-hop reasoning.
  3. 3.
    Future efforts will focus on evolving our instance-specific circuit optimizations into a generalized mitigation toolkit, supported by a comprehensive evaluation framework. Key improvements will target the precision of automated overshadowed knowledge identification and the broader efficacy of circuit-based interventions. Ultimately, we aim to develop PhantomCircuit as a robust platform for both in-depth analysis and effective, generalizable overshadowing mitigation.
Appendix A More Related Work
A.1 Large Language Models
First proposed by Brown et al. (2020), Transformer-based auto-regressive LLMs have demonstrated strong performance across a variety of NLP tasks, including question answering Yue (2025), in-context learning Dong et al. (2022), and analogical reasoning Webb et al. (2023). Pretrained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources. As a result, models such as InternLM2.5 Cai et al. (2024), Qwen2.5 Team (2024), and LLaMA3.3 Grattafiori et al. (2024) have shown excellent performance on world knowledge benchmarks Suzgun et al. (2022). Therefore, over the past year, LLMs have demonstrated remarkable capabilities in understanding-related tasks across various fields Kim et al. (2024); Nam et al. (2024); Yan et al. (2024c); Yan and Lee (2024); Yan et al. (2024b); Dong et al. (2025); Su et al. (2025).
Recently, there has been a growing trend toward enhancing LLMs’ reasoning capabilities on complex tasks Guo et al. (2025); Jaech et al. (2024) by generating long Chain-of-Thoughts (CoTs), with reinforcement learning (RL) emerging as an effective tool to encourage this behavior Li et al. (2025); Trung et al. (2024). Recently, there have also been efforts to explore collaboration between LLMs to enhance their reasoning abilities Zhang et al. (2025b); Putta et al. (2024); Masterman et al. (2024); Yan et al. (2025b); Chu et al. (2025).
Despite these advancements, existing LLMs still suffer from factual hallucinations in practice Pan et al. (2025); Asgari et al. (2025), with knowledge overshadowing identified as a primary contributing factor Zhang et al. (2024b). While existing interoperability works make great efforts on the mechanism of LLM training and generating Zhao et al. (2024), most of them solely focus on isolated model versions like GPT2 Wang et al. (2023) and LLaMA2 Wendler et al. (2024); Tang et al. (2024).
In this paper, we utilize the Pythia suite Biderman et al. (2023) to investigate the evolution and underlying mechanisms of knowledge overshadowing across models of varying sizes: 70M, 410M, 1.4B, and 2.8B parameters. Sharing a unified architecture, this model suite eliminates design variability, thereby providing clearer and more reliable insights into the scaling behavior of the knowledge overshadowing phenomenon in LLMs.
Appendix B PhantomCircuit Details
B.1 Circuit Construction
Knowledge circuit is as a sparse computational subgraph within the LLMs. The construction of such a circuit involves identifying and retaining the most influential components (nodes, including MLPs and attention heads) and connections (edges) while pruning less critical ones.
We adapted the optimized circuit construction method provided by Yao et al. (2024). The process begins by representing the LLM as a directed acyclic graph (DAG), $G=(V,E)$, where $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, and $E$ represents the information flow between these components. The goal is to identify a subgraph $C\subseteq G$ that is critical for recognize the key component of a given input prompt, particularly in knowledge overshadowing, is $X_{dom}$ and $X_{sub}$, the difference between $P_{dom}$ and $P_{sub}$.
The adapted construction method is similar to edge attribution patching(EAP) Conmy et al. (2023), which involves:
  1. 1.
    Paired Inputs: For a given background $X_{bg}$, we create two primary input prompts: $P_{dom}=(X_{bg},X_{dom})$ and $P_{sub}=(X_{bg},X_{sub})$. We also consider a "corrupted" version of $P_{sub}$, which could be $P_{dom}$ itself or another prompt designed to elicit $Y_{dom}$. Let’s denote the "clean" input as $P_{clean}$ (typically $P_{sub}$) and the "corrupted" input as $P_{corr}$ (designed to lead to $Y_{dom}$).
  2. 2.
    Activation Difference Calculation: We run both $P_{clean}$ and $P_{corr}$ through the model. For each node $v\in V$ that is a potential parent in an edge, we record its output activation. The difference in activations between the clean and corrupted runs for a node $v_{p}$ (parent) is denoted as $\Delta A(v_{p})=A_{clean}(v_{p})-A_{corr}(v_{p})$.
  3. 3.
    Edge Scoring via Gradient-based Attribution: To score an edge $e=(v_{p},v_{c})$ (from parent $v_{p}$ to child $v_{c}$), we focus on how patching the activation from $v_{p}$ (i.e., using $A_{clean}(v_{p})$ instead of $A_{corr}(v_{p})$ when $P_{corr}$ is the main input) affects a chosen metric $\mathcal{M}$. This metric $\mathcal{M}$ is designed to measure the model’s tendency towards generating $Y_{sub}$ versus $Y_{dom}$ when the input is $P_{sub}$. A common choice for $\mathcal{M}$ could be the logit difference between $Y_{sub}$ and $Y_{dom}$ at the final layer, or a metric related to our Relative Overshadowing rate ($RO$).
    The score $S(e)$ for an edge $e$ can be approximated by the product of the activation difference from its parent node and the gradient of the metric $\mathcal{M}$ with respect to the input of its child node, when the child node receives the "clean" activation from the parent while other inputs are "corrupted":
    $S(e)\approx\mathbb{E}_{P_{sub}}\left[\Delta A(v_{p})\cdot\frac{\partial\mathcal{M}(Y_{target}|P_{sub})}{\partial A_{input}(v_{c})}\right]$
    where $Y_{target}$ is ideally $Y_{sub}$. The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in our evaluation set $Z_{sub}$. In practice, methods like Integrated Gradients (IG) are often used to refine this attribution by integrating gradients along a path from a baseline (corrupted) input to the actual (clean) input.
  4. 4.
    Circuit Pruning: Based on the calculated scores $S(e)$, edges with scores below a certain threshold $\tau$, or alternatively, edges outside the top-N highest scores, are pruned from the graph $G$. The remaining nodes and edges form the knowledge circuit $C_{sub}$.
    $C_{sub}=(V_{sub},E_{sub})$
    where $E_{sub}=\{e\in E\mid|S(e)|\geq\tau\}$ (or top-N criterion) and $V_{sub}$ consists of nodes connected by edges in $E_{sub}$.
    This constructed circuit $C_{sub}$ is then analyzed to understand how dominant knowledge $K_{dom}$ might overshadow $K_{sub}$ by examining the attentional features and information flow within it, especially when processing $P_{sub}$.
B.2 Automated Component Identification for Recovery
Identifying the Overshadowed Component $X_{sub}^{*}$. A critical precursor to effective circuit-based recovery is the precise identification of the specific component $X_{sub}^{*}$ within the subordinate prompt $P_{sub}$ that is being overshadowed. This is achieved by adapting the Relative Pointwise Mutual Information (R-PMI) based methodology from Zhang et al. (2025a, 2024b). The process involves:
Iteratively generating contrastive prompts $P^{\prime}_{sub}$ by deleting each candidate token $X^{\prime}_{sub}$ (a potential overshadowed component) from the original $P_{sub}$.
For each pair $(P_{sub},P^{\prime}_{sub})$, calculating the R-PMI for tokens $y_{i}$ in the intersection of their top-$k$ next-token candidate sets, $V_{top}(P_{sub})\cap V_{top}(P^{\prime}_{sub})$, using
$R\textnormal{-}PMI(y_{i};P_{sub},P^{\prime}_{sub})={}$ $\log P(y_{i}|P_{sub})$
$-\log P(y_{i}|P^{\prime}_{sub}).$
Summing only the negative R-PMI values to obtain
$S_{R\textnormal{-}PMI^{-}}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI^{-}}$ is identified as the primary overshadowed component, $X_{sub}^{*}$. This selection is based on the rationale that removing the true $X_{sub}^{*}$ most strongly exposes the model’s bias towards outputs favored by the dominant knowledge pattern.
Identifying Target Subordinate Output $Y_{sub}$. The intended subordinate output $Y_{sub}$ is identified by assessing which token from $V_{top}(P_{sub})$ (the top-k candidates for the original prompt $P_{sub}=(X_{bg},X_{sub})$) exhibits the most significant improvement when the overshadowing influence of background knowledge ($X_{bg}$) or other non-subordinate components is mitigated. Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ (or other identified non-subordinate elements that contribute to the $X_{bg}\leftrightarrow Y_{dom}$ association) within $P_{sub}$. $Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement (or largest increase in log probability) in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$. This rank elevation signifies the “unmasking” of the true subordinate answer as the dominant, overshadowing associations are weakened.
Identifying Dominant Output $Y_{dom}$. The dominant output $Y_{dom}$ is identified as the token that maintains the highest average rank across all contrastive prompts $P^{\prime}_{sub}$ generated by deleting different candidate tokens $X^{\prime}_{sub}$ from $P_{sub}$. This token represents the model’s most consistent, default output tendency when specific subordinate cues are variably weakened, likely reflecting the pervasive influence of dominant knowledge associated with the background $X_{bg}$.
[…]
Some more circuit-based overshadowing recovery cases are shown in Table 1.
Table 1: Circuit-based overshadowing recovery cases
{tblr}
[…]
Rank 1: Logit: 15.42 Prob: 14.98% Token: |Time| Rank 1: Logit: 19.75 Prob: 20.42% Token: |Tem|
Rank 2: Logit: 14.21 Prob: 4.47% Token: |Sp| Rank 2: Logit: 17.34 Prob: 1.84% Token: |Filter|
Rank 3: Logit: 14.05 Prob: 3.83% Token: |T| Rank 3: Logit: 16.40 Prob: 0.72% Token: |E|
Rank 4: Logit: 13.84 Prob: 3.10% Token: |Custom| Rank 4: Logit: 16.29 Prob: 0.64% Token: |Sp|
[…]
Rank 1: Logit: 16.89 Prob: 17.65% Token: |E| Rank 1: Logit: 16.47 Prob: 10.39% Token: |Ps|
Rank 2: Logit: 15.35 Prob: 3.76% Token: |Ps| Rank 2: Logit: 15.47 Prob: 3.84% Token: |Wave|
Rank 3: Logit: 15.29 Prob: 3.57% Token: |St| Rank 3: Logit: 15.31 Prob: 3.28% Token: |Energy|
Rank 4: Logit: 15.15 Prob: 3.09% Token: |emitter| Rank 4: Logit: 14.89 Prob: 2.15% Token: |emitter|
[…]
Rank 1: Logit: 48.68 Prob: 33.09% Token: |St| Rank 1: Logit: 44.01 Prob: 27.96% Token: |Transfer|
Rank 2: Logit: 46.66 Prob: 4.38% Token: |Ada| Rank 2: Logit: 42.38 Prob: 5.48% Token: |St|
Rank 3: Logit: 46.37 Prob: 3.28% Token: |Flow| Rank 3: Logit: 42.24 Prob: 4.75% Token: |Energy|
Rank 4: Logit: 46.06 Prob: 2.41% Token: |Emit| Rank 4: Logit: 41.64 Prob: 2.61% Token: |Mi|
[…]
Rank 1: Logit: 38.34 Prob: 6.67% Token: |St| Rank 1: Logit: 34.24 Prob: 35.19% Token: |Field|
Rank 2: Logit: 35.87 Prob: 0.56% Token: |Fl| Rank 2: Logit: 31.41 Prob: 2.08% Token: |Output|
Rank 3: Logit: 35.82 Prob: 0.54% Token: |Trans| Rank 3: Logit: 30.97 Prob: 1.34% Token: |Har|
Rank 4: Logit: 33.51 Prob: 0.05% Token: |Un| Rank 4: Logit: 30.84 Prob: 1.18% Token: |Fl|
Appendix C Dataset Details
C.1 Detailed Synthetic Dataset Construction
The synthetic dataset was constructed through the following steps to ensure controlled conditions for analyzing knowledge overshadowing dynamics:
Fixing text lengths. For all generated data instances, consistent token lengths are maintained. The background knowledge ($X_{bg}$) was set to a length of 4 tokens. All other core components, namely the dominant knowledge entity ($X_{dom}$), subordinate knowledge entity ($X_{sub}$), dominant output ($Y_{dom}$), and subordinate output ($Y_{sub}$), are each set to a length of 1 token.
Dataset generation for specific D and P Combinations. For each defined combination of total dataset size (D) and knowledge popularity (P), the dataset was built as follows: The dataset comprises multiple distinct groups of knowledge instances. Each group consists of P+1 knowledge prompts: a set of P dominant knowledge prompts $\{P_{dom}^{1},P_{dom}^{2},\dots,P_{dom}^{P}\}$ and one subordinate knowledge prompt $P_{sub}$. Within each group:
  • For the P dominant prompts, the actual dominant knowledge entities $\{X_{dom}^{1},X_{dom}^{2},\dots,X_{dom}^{P}\}$ are all unique. However, they all share the same background knowledge component ($X_{bg}$) and are associated with the same dominant output ($Y_{dom\_g}$). Thus, each $P_{dom}^{i}=(X_{bg},X_{dom}^{i})$ is paired with $Y_{dom}^{g}$.
  • The single subordinate prompt $P_{sub}=(X_{bg},X_{sub})$ uses the same background knowledge $X_{bg}$ as the dominant prompts in that group. However, its subordinate knowledge entity $X_{sub}$ is distinct from all $X_{dom}^{i}$ entities in that group, and its corresponding output $Y_{sub}$ is distinct from any $Y_{dom}$ in group.
This structure creates a group: $\{(P_{dom}^{1},Y_{dom}^{1}),\dots,((P_{dom}^{P},Y_{dom}^{P}),(P_{sub},Y_{sub})\}$.
Multiple such groups are generated. All tokens for $X_{bg},X_{dom}^{i},X_{sub},Y_{dom},Y_{sub}$ within each group, and across different groups, are randomly sampled from the Pythia tokenizer vocabulary, ensuring no overlap between the core entities of different groups. This process was repeated until the total number of tokens in the dataset reached the target size D.
Cases illustration. We illustrate some groups for P=5 dataset in Table 2. We directly show token id.
Table 2: Illustrative examples from the synthetic dataset (P=5). Each data entry is a row, with fine lines separating entries within a group. Token IDs are shown.
{tblr}
[…]
$[10030,16936,1050,10565]$ $24327$ $20730$
$[10030,16936,1050,10565]$ $4619$ $20730$
$[10030,16936,1050,10565]$ $5137$ $20730$
$[10030,16936,1050,10565]$ $785$ $20730$
$[10030,16936,1050,10565]$ $18941$ $3519$
[…]
$[17026,8837,3802,28741]$ $3530$ $1077$
$[17026,8837,3802,28741]$ $11948$ $1077$
$[17026,8837,3802,28741]$ $2028$ $1077$
$[17026,8837,3802,28741]$ $9389$ $1077$
$[17026,8837,3802,28741]$ $25814$ $5374$
[…]
$[18131,14501,21161,311]$ $778$ $7790$
$[18131,14501,21161,311]$ $18762$ $7790$
$[18131,14501,21161,311]$ $28591$ $7790$
$[18131,14501,21161,311]$ $28981$ $7790$
$[18131,14501,21161,311]$ $8447$ $6129$
C.2 Finetuning dataset
For the finetuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge. This generated data subsequently underwent manual review to identify and remove any instances that are overly repetitive or semantically too similar, ensuring a degree of diversity, resulted in D = 1M.
A key distinction from the synthetic dataset construction is that we did not strictly control token lengths for each component in this dataset. Instead of randomly sampled token IDs, the finetuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge. The underlying pattern of dominant and subordinate knowledge construction, however, mirrors that of the synthetic dataset.
As an example of this dataset, we set the knowledge popularity P=5. Some illustrative cases from the dataset are shown in Table 3.
Table 3: Illustrative examples from the Finetuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
{tblr}
[…]
Analysis of the Starfire Crystal Engine reveals primary energy output idle cycle Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output phase synchronicity Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output null gravity Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output core temperature Pure Nova
Analysis of the Starfire Crystal Engine reveals primary energy output cold ignition Quantum Flux
[…]
Observation of the Chronos Orb’s temporal distortion signature indicates near Leyline Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates while attuned Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates during solstice Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates etherium interaction Stable Chrono-displacement
Observation of the Chronos Orb’s temporal distortion signature indicates sudden impact Temporal Anomaly
[…]
Study of the Xylosian biology reveals primary nutrient absorption method after sonic bath Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method during digestion Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method high pressure Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method thermal vent Molecular Assimilation
Study of the Xylosian biology reveals primary nutrient absorption method xenoflora consumption Crystalline Excretion
(a) The $\mathcal{RO}$ during training phase of the different finetuning dataset size (D).
(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on finetuning dataset.
Figure 6: The dynamics analysis of knowledge overshadowing in finetuning dataset.
Appendix D More Dynamics Analysis on Finetuning Dataset
In addition to validating the efficacy of our circuit-based overshadowing recovery method, the finetuning dataset serves a dual purpose. We also leverage it to empirically verify our conclusions regarding the training dynamics of knowledge overshadowing, specifically concerning the impact of Dataset Size (D). Consistent with our dynamic analysis findings, we investigate whether a larger D indeed correlates with a slower recovery rate from knowledge overshadowing and a prolonged duration of the hallucination effect. To this end, we conduct experiments on the finetuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens. The results, as depicted in Figure 6(a), corroborate this relationship. Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the finetuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$). As shown in Figure 6(b), the observations again support the hypothesis that insufficient optimization of subordinate knowledge contributes to the persistence of knowledge overshadowing.
It is noteworthy that distinct behaviors are observed when comparing the finetuning dataset to the synthetic dataset. Firstly, the recovery from overshadowing on the finetuning dataset is generally slower than on the synthetic dataset for same D. This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the finetuning data, which presents a more challenging learning task.
Secondly, we observe that the finetuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise. This is because finetuning commences from a pretrained model, which has already moved beyond the initial epochs of chaotic, random predictions. Consequently, the model can very rapidly generalize strong association patterns present in the finetuning data. Moreover, the diverse and varied forms of data within the finetuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences. This inherent data diversity can help preemptively mitigate or even eliminate the early onset stage of knowledge overshadowing that might otherwise be observed.